OpenStudy (anonymous):

H_E_L_P_!_!_! Suppose y varies directly with x, and y = 24 when x = 8. WHat is the value of y when x = 10 ? A) 10 B) 30 C) 80 D) 240

OpenStudy (anonymous):

What's K ?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

THANK YOU !

OpenStudy (anonymous):

8/24=1/3=10/30

OpenStudy (anonymous):

proportions